yomomalikebracobama:

Solve for x: 1/3 (2x − 8) = 4. (1 point) a 2 b 6 c 10 d 10/3

yomomalikebracobama:

so distribute

QuestionCoveBot:

Same as last one. 1/3 gets distributed to 2x and 8.

yomomalikebracobama:

how tho /idk how to distribute fraction

QuestionCoveBot:

I forgot how, but I'm not gonna bother looking. @mxddi3 ~

yomomalikebracobama:

@mxddi3 ~

yomomalikebracobama:

i dont think she/he is on or can help

QuestionCoveBot:

Maddie will be here when she can.

yomomalikebracobama:

do we do new question ?

QuestionCoveBot:

Not yet. I'll be busy soon, so Maddie might take over then.

yomomalikebracobama:

o ok lol i only have a hour left ima have to work after school work time ima jump off a bridge

yomomalikebracobama:

(not actually )

yomomalikebracobama:

hi sailor

yomomalikebracobama:

@astrid1

Sailor:

One third is distributed like @questioncovebot said, so you should get \[2/3x -8/3=4\]

Sailor:

Then you would subtract -8/3 from both sides

yomomalikebracobama:

ik buthow do iget that how do idistribute a fraction

yomomalikebracobama:

im no goood with fractions

Sailor:

Ok you distribute a fraction by multiplying it with whatever is in the parenthesis

yomomalikebracobama:

o ok

yomomalikebracobama:

so u make whole numbers like this say like 9/1 * 1/3

Sailor:

 Kinda like this

yomomalikebracobama:

/?

yomomalikebracobama:

ooo ok

Sailor:

When distributing a fraction it's probably best to have a calculator though

yomomalikebracobama:

yea i would think so lol

yomomalikebracobama:

so lets get the answer whats the nest step

Sailor:

You would subtract -8/3 from both sides and get \[2/3x=20/3\]

Sailor:

If you divide by both sides, you should get x=10

yomomalikebracobama:

ok thanks ima have alll my questionsdisplayed cuz idk how to do any of them

Sailor:

Ok. Glad to help

yomomalikebracobama:

how do i close it

yomomalikebracobama:

lol